OpenStudy (anonymous):

ln(x^2)=9 has two solutions x1

OpenStudy (anonymous):

x^2 = e^9 ->x = +/-sqrt( e^9)

OpenStudy (anonymous):

that didn't work....

OpenStudy (anonymous):

Correct answer.. x^2 = e^9 x=+e^3 or x=-e^3

OpenStudy (anonymous):

still not accepting those answers.... haha I'm not sure at all how to do this

OpenStudy (anonymous):

Sorry man , it is x^2 = e^9 x=+e^4.5 or x=-e^4.5

OpenStudy (anonymous):

thats still not right... haha this problem is crazy

OpenStudy (anonymous):